OpenStudy (anonymous):

another question @jim_thompson5910 @billj5

OpenStudy (anonymous):

I have to find the average value for this:

OpenStudy (anonymous):

\[\int\limits_{0}^{-3}\frac{ 4 }{ (2x+6)^{2} }\]

jimthompson5910 (jim_thompson5910):

you'll use u-substitution here

OpenStudy (anonymous):

since b is less then a then i have to switch them and do \[-\int\limits_{-3}^{0}\]

jimthompson5910 (jim_thompson5910):

yes, correct

OpenStudy (anonymous):

but then for my 1/b-a what do i use for b and what do i use for a?

jimthompson5910 (jim_thompson5910):

I think it doesn't matter really because the "b-a" portion should be positive anyway |b-a| = |a-b|

OpenStudy (anonymous):

I really can't figure this out, can you help me? @jim_thompson5910

OpenStudy (anonymous):

\[\int\limits_{0}^{-3}\frac{ 4 }{ (2x+6)^2 }\] and I gave to find the avg value and mean value

jimthompson5910 (jim_thompson5910):

did you try u-substitution?

OpenStudy (anonymous):

my teacher never told us to do that for this problem

jimthompson5910 (jim_thompson5910):

ok so let's replace 2x+6 with some simpler variable, often that's the letter u

jimthompson5910 (jim_thompson5910):

let u = 2x+6 if u = 2x+6, then du/dx = 2, agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

that means du/dx = 2 du = 2*dx I multiplied both sides by dx

jimthompson5910 (jim_thompson5910):

then multiply both sides by 2 du = 2*dx 2du = 4dx

jimthompson5910 (jim_thompson5910):